Shift summary for owl 8/21/99

Jason Hamm (jhamm@physics.arizona.edu)
Sat, 21 Aug 1999 04:59:46 -0700 (MST)


Shift Summary for owl shift, 08/21/99

On Shift: J. Hamm, T. Alexopolous

Hours of beam available at all intensities: 6.3
Hours of beam available at nominal intensity: 6.3
Hours of E832 data (special runs not included): 6.2
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Beamline magnet problems, stream 5 errors
========================================================================
Runs ended this shift:

none (14291 still in progress)

Tapes brought to FCC this shift: QK(S,C,K)294-296, QKB103
QKA013-014, QKLA13, QKSM12, QKCC13,
QKNL22-23, QKN024-025
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Mostly smooth running at ~6.8 E12
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
0:05 Start Beam_832 run 14291. No ktevana errors at
start of run!

14291 73 1:45 Unknown crate, stream 5. Kfrend, miss spills
73-74

14291 90 2:10 Get "RC10/REG_TOP out of limits" errors - scalar
screen claims RC10 is counting about a factor of
5 too low. Online histos look OK, though. Scalar
back to normal after 5 spills.

14291 2:15 Onsplit requests QKN026 on ktev1.

14291 2:45 Onsplit requests QKA015 on ktev3.

14291 147 3:25 Missing FFFF word, stream 5. Kfrend, miss this
spill. RC scalar errors reappear. Online histos
still look OK. Back to normal after 6 spills.

14291 151 3:30 DYC-53 full. Lose this spill.

14291 207 4:45 RC10 error again. Online histos still look OK.
back to normal after 6 spills.

14291 4:55 More stream 5 errors.

14291 5:05 Lose beam due to NM1U trip.

14291 248 5:40 Beam reappears, along with RC10 problem. Scalar
reads low 10 spills in a row, so page SET, who
thinks this is a bad cable/connection at patch
panel; since histos look ok, we can ignore.
Theo changes alarm limit for this scalar so we
don't get beeped at continuously.

14291 272 6:10 More stream 5 errors; lose spills 272-276 while
kfrending, cycling repeater, etc.

14291 291 6:35 No beam through spill 295.
========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift